# Topological phase transition through tunable nearest-neighbor interactions in a one-dimensional lattice

### v1

@@ III.3 Away from half-filling
After studying the ground state phase diagram in both the regimes of interactions at half-filling, we now extend our analysis to explore the system across all densities. For this purpose, we again focus on the same cut through the phase diagram that corresponds to the $V_{1}=-V_{2}=V$ line (gray dashed line in Fig. 2 (a)) as mentioned earlier. The phase diagram of the system in the $V-\mu$ plane is depicted in Fig. 12. The gapped phases at $\rho=1/2$, namely the BO and the CDW-II phases, appear in the regions with $V>0$ and $V<0$ and are separated by a vertical dashed line. At $V=0$ the gap vanishes and it corresponds to the topological phase transition point between the BOπ and BO0 phases at half-filling as discussed earlier. The gray regions on the upper and lower sides of the gapped phases are the gapless phases which appear for densities away from half-filling. The upper and lower white regions are the completely filled and completely empty (vacuum) states with $\rho=1$ and $\rho=0$ respectively. The boundaries of these phases are determined from the change in the chemical potential $\mu$ as the particle density $\rho$ of the system is defined. For this purpose, we analyze the features from the $\rho$ vs $\mu$ plot which are shown in Figs. 13 (a) and (b) for two exemplary values of $V=-3.0$ and $V=-5.5$ respectively.
Figure 12: The phase diagram of the model in $V-\mu$ plane by taking $V_{1}=-V_{2}=V$ under open boundary condition. Dashed lines indicate the presence of edge states. The white regions in the middle enclosed by gray regions correspond to gapped phases at half-filling, while the white regions at $\rho=0$ and $\rho=1$ represent the empty and fully occupied states, respectively. Gray-shaded regions denote gapless phases, which are further classified based on excitation type: regions with single-particle excitations are identified as the SF phase, while regions with two-particle excitations correspond to the PSF phase. Red solid lines mark the boundary between the SF and PSF phases.
Here we plot the bulk and the edge densities defined as
@@
using red solid lines and blue dashed lines respectively. The large plateaus in $\rho_{b}$ (red solid lines) in both the figures at $\rho=1/2$ is due to the gapped phases, and the shoulders across the plateaus correspond to the gapless phases. The length of the plateau at $\rho=1/2$ determines the gap at this filling, and the end points of the plateau determine the boundaries of the gapped phase appearing at $\rho=1/2$ corresponding to the particular cut in the phase diagram shown in Fig. 12. Interestingly, there are two degenerate mid-gap zero energy modes appearing in the phase diagram starting from $V=0$ to $V\sim-4.3$ surpassing which these states become energetic, and these are indicated by the dashed lines at the center of the phase diagram. These are the signatures of edge states which have already been discussed in the previous section; they appear here as a consequence of the bulk topology. The degeneracy persists in the topological BO phase which then becomes non-degenerate because of the transition into the trivial CDW-II phase.
This signature of edge states can be captured from the $\rho-\mu$ plot shown in Fig. 13. The jump in $\rho_{e}$ by 1 when the bulk is gapped in Fig. 13 (a) at $\mu=0$ is an indication of the presence of two zero energy edge states in this parameter regime. Similarly, the jump in $\rho_{e}$ by 0.5 in Fig. 13 (b) while the bulk is gapped at finite $\mu$ represents the presence of finite-energy edge states.
@@
Figure 13: The figure depicts the bulk and edge densities represented by $\rho_{b}$ (red solid lines) and $\rho_{e}$ (blue dashed lines) respectively as a function of $\mu$, for two different cuts through the phase diagram shown in Fig. 12: (a) $V=-3.0$ and (b) $V=-5.5$ for a system size of $L=200$. This shows the nature of bulk phases (gapped or gapless) when the edge states are being filled in different parameter regimes. The inset shows the number of particles filled $N$ as a function of chemical potential $\mu$. The single-particle and two-particle jumps in the insets of (a) and (b) represent the existence of SF and PSF phases respectively.
Figure 14: The SF correlation function $\Gamma\text{(}r)$ and the PSF correlation function $\Gamma^{\text{2}}(r)$ plotted as a function of the distance ($r=|i-j|$) on a log-log scale at $\rho=0.25$ for $V=5.5$ for a system with size $L=200$. To avoid the boundary effects we only consider the lattice sites from $L/4$ to $43L/$ 4.
Figure 15: The figure shows the number of particles filled as a function of the chemical potential $\mu$ for $V=5.5$ under PBC for a system size of $L=40$. The inset shows the zoomed region of the red curve which is indicated by an arrow. The two-particle jump is clear from the inset which is a signature of the PSF phase.
Figure 16: Phase diagram of the model in the $V-\mu$ plane by taking $V_{1}=-V_{2}=V$ with periodic boundary condition. All the phase boundaries are extrapolated by taking a maximum system size of $L=60$. All the regions have the usual meaning as mentioned in the phase diagram shown in Fig. 12. The red dashed lines mark the boundary between the SF and the PSF phases. The rectangular region marked in blue in the upper right part of the diagram is enlarged in the inset to show that the $\rho=0$ and $\rho=1/2$ regions are separated by a finite gray region which carries PSF order.
@@
We now examine the phases away from commensurate densities. As evident from the phase diagram [Fig. 12] as well as from the $\rho-\mu$ curve [Fig. 13], the regions away from half-filling correspond to gapless phases. In some regions, the excitations are of single-particle type representing an SF phase, whereas in some other regions, the excitations are pair excitations representing a pair superfluid (PSF) phase. The PSF phase arises due to strong attractive interactions that favor the formation of bound nearest-neighbor pairs. To distinguish between the SF and PSF phases, we plot the particle number $N$ as a function of the chemical potential $\mu$ in the inset of Fig. 13, focusing on the region marked by the black rectangle for a system size of $L=200$ sites. The single-particle jump in $N$ with increase in $\mu$ for $V=-3.0$ (inset in Fig. 13 (a)) shows the presence of an SF phase, whereas the two-particle jump for $V=-5.5$ (inset in Fig. 13 (b)) shows the existence of the PSF phase. These features can also be understood from analytical arguments discussed in Appendix B. The jumps in $\rho_{b}$ in steps of $1$ occurring when $\mu$ is slightly more than $-2.0$ (Fig. 13(a)) is due to single particles appearing in the bulk when $\mu$ increases above $-2.0$ as discussed in Appendix B (see Eq. (36)). The jumps in $\rho_{b}$ in steps of $2$ occurring when $\mu$ increases above $-2.94$ (Fig. 13(b)) is due to the appearance of two-particle bound states in the bulk. If we substitute $V=-5.5$ in Eq. (51), it gives a threshold equal to about $-2.95$ which matches well with the numerical results.
@@
The SF and PSF boundaries are separated by the red lines in the phase diagram shown in Fig. 12. Note that a similar SF-PSF transition line is observed for the $V>0$ case; however, it is not seen from the $\rho$-$\mu$ curve obtained with OBC because of the choice of the boundary condition and the repulsive interaction in the first and the last bonds. More details can be found in Appendix B. Nevertheless, the PSF nature can still be verified by examining the behavior of the PSF correlation function which is given by
@@
The algebraic decay of the PSF correlation function $\Gamma^{2}(r)$ (green circles) in Fig. 14 for $V=5.5$ shows the existence of a PSF phase in this parameter regime. For comparison, we also plot the SF correlation function $\Gamma(r)$ (blue squares) defined in Eq. 10 which shows an exponential decay confirming the existence of the PSF order in the system.
@@
To address this issue arising because of OBC, periodic boundary conditions (PBC) can be employed, allowing for a clearer examination of the $\rho$-$\mu$ curve. Hence, we analyze the $\rho-\mu$ curve at $V=5.5$ by employing PBC for a system size of $L=40$ and show the result in Fig. 15. The presence of a two-particle jump away from half-filling under PBC confirms the existence of the PSF phase in the bulk. The jumps occur due to the appearance of two-particle bound states in the bulk as discussed in Appendix B (see Eq. (51)). In contrast, under OBC, a single-particle density jump is observed even within the PSF region. Owing to this specific artifact introduced by the boundary condition, the phase diagram in Fig. 12 is not symmetric about $V=0$ although the bulk physics remains the same on the two sides. We provide a qualitative understanding of this picture in Appendix B; see the discussion around Eqs. (57) and (58) of hole or particle end modes which appear for $V>2.0$.
@@
To resolve this issue numerically, we once again compute the bulk phase diagram by implementing the PBC; this is presented in Fig. 16. The maximum system size in this case has been taken to be $L=60$, and all the phase boundaries are extrapolated to the thermodynamic limit. The mid-gap edge modes are now absent in the $V<0$ regime, as the implementation of PBC eliminates edge states. Now the phase diagram exhibits complete reflection symmetry about $V=0$. Additionally, there are four red lines in the phase diagram that mark the boundaries between the SF and PSF phases. One can clearly see that at higher interaction strengths, the boundary region between the vacuum state ($\rho=0$) and the half-filled state ($\rho=1/2$) (or between the fully occupied state ($\rho=1$) and the half-filled state) appears narrow and seemingly merge to a single line. However, a finite intermediate gray region remains, consisting of two-particle excitations, which defines the PSF phase. To illustrate this more clearly, the inset shown in Fig. 16 provides a zoomed-in view of the blue rectangular region, confirming the existence of a gapless phase between the two gapped phases.
@@
Figure 17: Array of dipoles representing the model under consideration. The angles $\theta$ and $\phi$ control the strength and signs of the interactions $V_{1}$ and $V_{2}$ respectively.
@@ III.4 Experimental scheme
In this subsection, we briefly discuss the possible experimental realization of the model with tunable NN interactions. As already mentioned in the introduction, the NN interactions have been successfully realized in various quantum simulators. Here, we show that a properly arranged array of dipolar atoms can be a suitable choice to achieve the tunability of the potentials $V_{1}$ and $V_{2}$ that is required to observe the topological phase transition. In Fig. 17, we show that the atoms arranged in the zig-zag pattern can selectively allow the atoms to hop between the NN sites while the hopping to the next-nearest-neighbor sites can be avoided by increasing the angle $\alpha$. We assume that the dipoles can be oriented in the $XY$-plane. The dipole-dipole interaction between the two dipoles is given by
@@
where $\Phi$ is the angle made by the dipoles with the line joining the two dipoles Giovanazzi et al. (2002). Consequently, the nearest-neighbor interaction strengths are given by
@@
where $\theta$ and $\phi$ are the angles specified in Fig. 17. Hence to tune the interaction between the NN dipoles as required by the model under consideration, $V_{1}$ and $V_{2}$ can be tuned by uniformly rotating all the dipoles. Here, we set $\alpha=2\theta_{m}$, where $\theta_{m}$ is the ”magic angle” — the angle at which the dipole-dipole interaction strength vanishes which is given by $\text{cos}^{-1}(\frac{1}{\sqrt{3}})$. In this configuration, the angle $\phi$ is given by $\pi+\theta-2\theta_{m}$. By externally controlling the dipole orientation in the $XY$-plane, it is possible to adjust $\theta$, which in turn modifies $\phi$. It then becomes possible to simultaneously set the angles $\theta$ and $\phi$ to be $\theta_{m}$ and $\pi-\theta_{m}$ which makes both the nearest-neighbor interaction strengths zero. Fig. 18 presents the variation of $V_{1}$ and $V_{2}$ as a function of $\theta$, assuming the prefactor in Eq. 27 to be unity. Initially, $V_{1}$ (red squares) is negative, while $V_{2}$ (green circles) is positive. As $\theta$ reaches $\theta_{m}$, both interactions vanish. Further increasing $\theta$ beyond this point reverses their signs. This controlled variation of interaction strengths enables the realization of topological phase transitions in the system.
Figure 18: Tunable NN interactions $V_{1}$ (red squares) and $V_{2}$ (green circles) as a function of $\theta$.
IV Conclusion

### v2

@@ III.3 Away from half-filling
After studying the ground state phase diagram in both the regimes of interactions at half-filling, we now extend our analysis to explore the system across all densities. For this purpose, we again focus on the same cut through the phase diagram that corresponds to the $V_{1}=-V_{2}=V$ line (gray dashed line in Fig. 2 (a)) as mentioned earlier. The phase diagram of the system in the $V-\mu$ plane is depicted in Fig. 14. The gapped phases at $\rho=1/2$, namely the BO and the CDW-II phases, appear in the regions with $V>0$ and $V<0$ and are separated by a vertical dotted line. At $V=0$ the gap vanishes and it corresponds to the topological phase transition point between the BOπ and BO0 phases at half-filling as discussed earlier. The gray regions on the upper and lower sides of the gapped phases are the gapless phases which appear for densities away from half-filling. The upper and lower white regions are the completely filled and completely empty (vacuum) states with $\rho=1$ and $\rho=0$ respectively. The boundaries of these phases are determined by following the analysis performed in Fig. 4. In this case however, we plot the bulk and the edge densities defined as
@@
in Figs. 15 (a) and (b) for two exemplary values of V=3.0V=-3.0 and V=5.5V=-5.5 as red solid lines and blue dashed lines, respectively.
Figure 14: The phase diagram of the model in $V-\mu$ plane by taking $V_{1}=-V_{2}=V$ under OBC. Dashed lines indicate the presence of edge states. The white regions in the middle enclosed by gray regions correspond to gapped phases at half-filling, while the white regions at $\rho=0$ and $\rho=1$ represent the empty and fully occupied states, respectively. Gray-shaded regions denote gapless phases, which are further classified based on excitation type: regions with single-particle excitations are identified as the SF phase, while regions with two-particle excitations correspond to the PSF phase. The solid magenta lines mark the boundaries between the SF and PSF phases. To clearly show this, we plot the zoomed-in portion demarcated by the blue rectangular box as shown in the inset.
@@
The large plateaus in $\rho_{b}$ (red solid lines) in both the figures at $\rho=1/2$ is due to the gapped phases, and the shoulders across the plateaus correspond to the gapless phases. The length of the plateau at $\rho=1/2$ determines the gap at this filling, and the end points of the plateau determine the boundaries of the gapped phase appearing at $\rho=1/2$ corresponding to the particular cut in the phase diagram shown in Fig. 14. Interestingly, there are two degenerate mid-gap zero energy modes appearing in the phase diagram starting from $V=0$ to $V\sim-4.3$ surpassing which these states become energetic, and these are indicated by the dashed lines at the center of the phase diagram. These are the signatures of edge states which have already been discussed in the previous section; they appear here as a consequence of the bulk topology. The degeneracy persists in the topological BO phase which then becomes non-degenerate because of the transition into the trivial CDW-II phase. This signature of edge states can be captured from the $\rho-\mu$ plot shown in Fig. 15. The jump in $\rho_{e}$ by 1 when the bulk is gapped in Fig. 15 (a) at $\mu=0$ is an indication of the presence of two zero energy edge states in this parameter regime. Similarly, the jump in $\rho_{e}$ by 0.5 in Fig. 15 (b) while the bulk is gapped at finite $\mu$ represents the presence of finite-energy edge states.
@@
Figure 15: The figure depicts the bulk and edge densities represented by $\rho_{b}$ (red solid lines) and $\rho_{e}$ (blue dashed lines) respectively as a function of $\mu$, for two different cuts through the phase diagram shown in Fig. 14: (a) V=3.0V=-3.0 and (b) V=5.5V=-5.5 for a system size of $L=200$. This shows the nature of bulk phases (gapped or gapless) when the edge states are being filled in different parameter regimes. The inset shows the number of particles filled (NN) as a function of chemical potential ($\mu$). The single-particle and two-particle jumps in the insets of (a) and (b) represent the existence of SF and PSF phases respectively.
@@
We now examine the phases away from commensurate densities. As evident from the phase diagram [Fig. 14] as well as from the $\rho-\mu$ curve [Fig. 15], the regions away from half-filling correspond to gapless phases. In some regions, the excitations are of single-particle type representing an SF phase, whereas in some other regions, the excitations are pair excitations representing a pair superfluid (PSF) phase. The PSF phase arises due to strong attractive interactions that favor the formation of bound nearest-neighbor pairs. To distinguish between the SF and PSF phases, we plot the particle number $N$ as a function of the chemical potential $\mu$ in the inset of Fig. 15, focusing on the region marked by the black rectangle for a system size of $L=200$ sites. The single-particle jump in $N$ with increase in $\mu$ for V=3.0V=-3.0 (inset in Fig. 15 (a)) shows the presence of an SF phase, whereas the two-particle jump for V=5.5V=-5.5 (inset in Fig. 15 (b)) shows the existence of the PSF phase. These features can also be understood from analytical arguments discussed in Appendix B. The jumps in $\rho_{b}$ in steps of $1$ occurring when $\mu$ is slightly more than 2.0-2.0 (Fig. 15(a)) is due to single particles appearing in the bulk when $\mu$ increases above 2.0-2.0 as discussed in Appendix B (see Eq. (37)). The jumps in $\rho_{b}$ in steps of $2$ occurring when $\mu$ increases above 2.94-2.94 (Fig. 15(b)) is due to the appearance of two-particle bound states in the bulk. If we substitute V=5.5V=-5.5 in Eq. (52), it gives a threshold equal to about 2.95-2.95 which matches well with the numerical results.
@@
The SF and PSF boundaries are separated by the magenta lines in the phase diagram shown in Fig. 14. Note that a similar SF-PSF transition line is observed for the $V>0$ case; however, it is not seen from the $\rho$-$\mu$ curve obtained with OBC because of the choice of the boundary condition and the repulsive interaction in the first and the last bonds. More details can be found in Appendix B. Nevertheless, the PSF nature can still be verified by examining the behavior of the PSF correlation function which is given by
@@
The algebraic decay of the PSF correlation function $\Gamma^{2}(r)$ (green circles) in Fig. 16 for $V=5.5$ at filling $\rho=0.25$ shows the existence of a PSF phase in this parameter regime. For comparison, we also plot the SF correlation function $\Gamma(r)$ (blue squares) defined in Eq. 9 which shows an exponential decay confirming the existence of the PSF order in the system.
Figure 16: The SF correlation function $\Gamma\text{(}r)$ and the PSF correlation function $\Gamma^{\text{2}}(r)$ plotted as a function of the distance (r=|ij|r=|i-j|) on a log-log scale at $\rho=0.25$ for $V=5.5$ for a system with size $L=200$. To avoid the boundary effects we only consider the lattice sites from $L/4$ to $3L/4$.
Figure 17: The figure shows the number of particles filled (NN) as a function of the chemical potential ($\mu$) for $V=5.5$ under PBC for a system size of $L=40$. The inset shows the zoomed-in region of the red curve which is enclosed within the black rectangle. The two-particle jump is clear from the inset which is a signature of the PSF phase.
@@
To address this issue arising because of OBC, the boundary condition can be changed to PBC which allows for a clearer examination of the $\rho-\mu$ curve. Hence, we analyze the $\rho-\mu$ curve at $V=5.5$ by employing PBC for a system size of $L=40$ and show the result in Fig. 17. The presence of a two-particle jump away from half-filling under PBC confirms the existence of the PSF phase in the bulk. The jumps occur due to the appearance of two-particle bound states in the bulk as discussed in Appendix B (see Eq. (52)). In contrast, under OBC, a single-particle density jump is observed even within the PSF region. Owing to this specific artifact introduced by the boundary condition, the phase diagram in Fig. 14 is not symmetric about $V=0$ although the bulk physics remains the same on the two sides. We provide a qualitative understanding of this picture in Appendix B; see the discussion around Eqs. (58) and (59) of hole or particle end modes which appear for $V>2.0$.
Figure 18: Phase diagram of the model in the $V-\mu$ plane by taking $V_{1}=-V_{2}=V$ with PBC. All the phase boundaries are extrapolated by taking a maximum system size of $L=60$. All the regions have the usual meaning as mentioned in the phase diagram shown in Fig. 14. The solid magenta lines mark the boundary between the SF and the PSF phases. To clearly show this, we plot the zoomed-in portion demarcated by the green rectangular box in the left inset. The rectangular region marked in blue in the upper right part of the diagram is enlarged in the right inset to show that the $\rho=1$ and $\rho=1/2$ regions are separated by a finite gray region that carries the PSF order.
@@
To resolve this issue numerically, we once again compute the bulk phase diagram by implementing the PBC; this is presented in Fig. 18. The maximum system size in this case has been taken to be $L=60$, and all the phase boundaries are extrapolated to the thermodynamic limit. The mid-gap edge modes are now absent in the $V<0$ regime, as the implementation of PBC eliminates the edge states. Now the phase diagram exhibits complete reflection symmetry about $V=0$. Additionally, there are four magenta lines in the phase diagram that mark the boundaries between the SF and PSF phases. One can clearly see that at higher interaction strengths, the boundary region between the vacuum state ($\rho=0$) and the half-filled state ($\rho=1/2$) (or between the fully occupied state ($\rho=1$) and the half-filled state) appears narrow and seemingly merge to a single line. However, a finite intermediate gray region remains, consisting of two-particle excitations, which defines the PSF phase as illustrated in the right inset of Fig. 18.
@@ III.4 Experimental scheme
Figure 19: Array of dipoles representing the model under consideration. The angles $\theta$ and $\phi$ control the strength and signs of the interactions $V_{1}$ and $V_{2}$, respectively. $\alpha$ decides the strength of the next-nearest-neighbour hopping strengths.
Figure 20: Tunable NN interactions $V_{1}$ (red squares) and $V_{2}$ (green circles) are plotted as a function of $\theta$. Here, the values of the interactions are scaled according to the prefactor $\frac{q^{2}}{4\pi\epsilon_{0}r^{3}}$ given in Eq. 26.
In this subsection, we briefly discuss the possible experimental realization of the model with tunable NN interactions. As already mentioned in the introduction, the NN interactions have been successfully realized in various quantum simulators. Here, we show that a properly arranged array of dipolar atoms loaded in an optical lattice can be a suitable platform to achieve the tunability of the potentials $V_{1}$ and $V_{2}$ that is required to observe the topological phase transition. In Fig. 19, we show that the atoms arranged in the zig-zag pattern can selectively allow the atoms to hop between the NN sites while the hopping to the next-nearest-neighbor sites can be avoided by increasing the angle $\alpha$. We assume that the dipoles can be oriented in the $XY$-plane. The dipole-dipole interaction between the two dipoles is given by
@@
where $\Phi$ is the angle made by the dipoles with the line joining the two dipoles Giovanazzi et al. (2002). Consequently, the NN interaction strengths are given by
@@
where $\theta$ and $\phi$ are the angles specified in Fig. 19. Hence to tune the interaction between the NN dipoles as required by the model under consideration, $V_{1}$ and $V_{2}$ can be tuned by uniformly rotating all the dipoles. Here, we set $\alpha=2\theta_{m}$, where $\theta_{m}$ is the ”magic angle” — the angle at which the dipole-dipole interaction strength vanishes which is given by $\text{cos}^{-1}(\frac{1}{\sqrt{3}})$. In this configuration, the angle $\phi$ is given by $2\theta_{m}-\theta$. By externally controlling the dipole orientation in the $XY$-plane, it is possible to adjust $\theta$, which in turn modifies $\phi$. It then becomes possible to simultaneously set both the angles $\theta$ and $\phi$ to be $\theta_{m}$ which makes both the NN interaction strengths zero. Fig. 20 presents the variation of $V_{1}$ and $V_{2}$ as a function of $\theta$, assuming the prefactor in Eq. 26 to be unity. Initially, $V_{1}$ (red squares) is negative, while $V_{2}$ (green circles) is positive. As $\theta$ reaches $\theta_{m}$, both interactions vanish. Further increasing $\theta$ beyond this point reverses their signs. This controlled variation of interaction strengths enables the realization of topological phase transitions in the system.
Figure 21: Phase diagram in the full $V_{1}$$V_{2}$ parameter space at half-filling, showing five distinct phases: bond order (BO), superfluid (SF), charge-density wave-I (CDW-I), charge-density wave-II (CDW-II), and phase separation (PS). White squares mark the BO to CDW-I phase boundary, white diamonds indicate the BO to CDW-II boundary, and the region enclosed by white circles corresponds to the PS phase. The red dashed line marks the SF phase.
[…]
IV Complete Phase Diagram at half-filling
In this section, we present the full phase diagram in the $V_{1}-V_{2}$ parameter space in Fig. 21, to provide a comprehensive overview of the ground state phases of the model. The second ($V_{1}<0,V_{2}>0$) and the fourth ($V_{1}>0,V_{2}<0$) quadrants have been analyzed in detail in the preceding section, while the first quadrant ($V_{1},V_{2}>0$) was previously studied in Ref. Mondal et al. (2022a); Nigam et al. (2025). It has already been shown that the first quadrant contains both the BO0 and BOπ phases, separated by a topological phase transition in the weak interaction regime (indicated by the red dashed line in Fig. 21). In addition to this, a charge-density wave phase (CDW-I), characterized by the occupation pattern of $\cdots 01010101\cdots$, appears within the region enclosed by white squares with increase in interactions.
To understand the scenario in the third quadrant ($V_{1},V_{2}<0$), we first examine the symmetric line which corresponds to $V_{1}=V_{2}$. Along this line, the model maps exactly to a spin-$1/2$ XXZXXZ chain, where $V_{2}(=V_{1})$ controls the strength of anisotropy and the $XX$ interaction strength is 2t-2t. The $XXZ$ model, solvable via the Bethe Ansatz Sutherland (2004), exhibits a gapless $XY$-ordered phase (equivalent to a SF phase in the bosonic language) for $2t<V_{2}<2t$ Cabra et al. (1998), as indicated by the red dashed line in Fig. 21. At $V_{2}=-2t$, a first-order quantum phase transition is known to occur from the $XY$ ordered phase to a ferromagnetic phase, which is also the phase-separated (PS) phase in the bosonic context. This PS phase features a clustered density profile of the form $11110000$, and the phase transition from the SF to the PS phase is signaled by a kink in the ground state energy $E$, indicating a non-analyticity in its derivative Justino and de Oliveira (2012). From our DMRG simulations, we explore this phase transition and the entire phase diagram in the $V_{1}-V_{2}$ plane corresponding to the third quadrant as depicted in Fig. 21. We obtain that the BO phases of the second and the fourth quadrants extend to the third quadrant and are separated by the topological phase transition line at $V_{1}=V_{2}$ (red dashed line) up to $V_{1}=V_{2}=2$. The CDW-II phases also extend to this quadrant at smaller values of interactions. Eventually, all these phases undergo transitions to the PS phase at the phase boundary denoted by circles. In the following, we characterize these phases and the transitions between them.
While we characterize the BO and the CDW-II phases by their respective order parameters discussed above, we identify the PS phase from the spatial distribution of the onsite density $\langle n_{i}\rangle$. In Fig. 22(a) we plot the $\langle n_{i}\rangle$ vs $i$ for $V_{1}=-2.0$ and $V_{2}=-4.0$ which clearly shows the accumulation of all the particles in a region of the lattice indicating a phase separation. To quantify the PS phase, we define an order parameter as
$\text{O}_{\text{PS}}=\frac{1}{L/2}\left|\sum_{i=1}^{L/2}\langle n_{i}\rangle-\sum_{i=L/2+1}^{L}\langle n_{i}\rangle\right|,$ (27)
[…]
In the limit of large $|V_{1}|$ (or large $|V_{2}|$), the boundary between one of the CDW-II phases and the PS phase can be found using second-order perturbation theory; this is presented in Appendix C. For instance, as $V_{1}\to$-$\infty$, we find that there is a first-order phase transition of a classical Ising model at a line given by
$V_{2}~=~\frac{2t^{2}}{V_{1}}.$ (28)
This asymptotically approaches $V_{2}=0$ from the negative side as $V_{1}\to$-$\infty$, as we see near the middle of the left edge of Fig. 21.
V Conclusion